### v1

# Path planning with moving obstacles using stochastic optimal control

Abstract
Navigating a collision-free, optimal path for a robot poses a perpetual challenge, particularly in the presence of moving objects such as humans. In this study, we formulate the problem of finding an optimal path as a stochastic optimal control problem. However, obtaining a solution to this problem is nontrivial. Therefore, we consider a simplified problem, which is more tractable. For this simplified formulation, we are able to solve the corresponding Bellman equation. However, the solution obtained from the simplified problem does not sufficiently address the original problem of interest. To address the full problem, we propose a numerical procedure where we solve an optimization problem at each sampling instant. The solution to the simplified problem is integrated into the online formulation as a final-state penalty. We illustrate the efficiency of the proposed method using a numerical example.
I INTRODUCTION
There are numerous applications of autonomous mobile robots that are discussed in the literature. In general, these applications are divided into two groups: indoor applications, such as delivering packages, cargo, and cleaning large buildings [1], and outdoor field robotics application [2]. Finding an optimal path for the robot to reach its destination is a crucial task in these applications. Path planning is one of the primary challenges that must be solved before mobile robots can autonomously navigate and explore complex environments [3]. Based on the nature of the environment that the robot is working in, the mobile path planning is divided into two classes, i.e. static environment and dynamic environment [1][4]. In the static environment, the obstacles are static and do not move. However, in the dynamic environment, the obstacles move during the planning task, while the robot is moving to reach its destination. The primary goal of path planning is to ensure safe, efficient, and collision-free navigation in both static and dynamic environments [5]. Path planning algorithms can be broadly categorized into several groups. The first group is classical algorithms, such as Dijkstra’s algorithm. The second group is heuristic-based approaches, e.g., the $A^{\star}$ algorithm. The third group is artificial intelligence-based approaches, such as reinforcement learning, fuzzy logic, and so on. The fourth group is hybrid-based approaches, which combine elements from all the above classes of path planning [1]. However, this is not the only way to classify path planning algorithms. In [4][6], and [7], other classifications for different approaches to path planning are also presented.
In this paper, we consider a case in which there is a robot and a stochastically moving obstacle in the environment. Also, we assume that there is a static target, and we want to control the robot to reach this target. We formulate the problem of finding an optimal path as an infinite horizon stochastic optimal control problem. Numerous deterministic optimal control problem formulations are available for path planning [8]; however, to the best of the author’s knowledge, the stochastic setting has not been considered. The associated stochastic Bellman equation can be solved using value iteration. Notably, for many problems, achieving an exact solution is simply unfeasible due to what Bellman properly termed the ‘curse of dimensionality,’ in his book [9, p. 94]. This problem refers to the exponential increase in the required computations as the problem size grows [10, p. 282]. In order to circumvent this problem, we proposed a method based on geometrical symmetry to reduce the dimension of the domain of the value function. Then we use fitted value iterations to find the value function. Fitted value iterations is a parametric approximation method in which we parameterize the value function as a linear regression such that it closely satisfies the Bellman equation [11, p. 211]. Fitted value iteration is a special case of approximate value iteration, dating back to 1959 [12]. We do not explicitly compute the policy, since it is defined in the original domain of the value function. Instead, we compute the control signal in real time for the given current values of the states.
We will compare our method with a receding horizon $A^{\star}$ algorithm. $A^{\star}$ is one of the most popular path finding algorithms [13]. In this paper, we consider a receding horizon version of this algorithm. We assume that the environment is static and find the path to reach the target using the $A^{\star}$ algorithm. At each stage, we take one step and discard the rest of the path generated by the $A^{\star}$ algorithm. One of the main drawbacks of this method is its large execution time [13].
Summarizing, the main contributions of this paper are:
  1. 1.
    We formulate an optimal path planning problem as a stochastic optimal control problem.
  2. 2.
    We show that the corresponding value function exhibits symmetry, enabling its representation with a reduced number of variables.
  3. 3.
    We use a numerical example to illustrate that the proposed framework can outperform the widely used $A^{\star}$ algorithm on the considered class of problems.
II Mathematical Model
Consider the problem where there is a robot, a stochastically moving obstacle, and a static target in a $2$-dimensional space. The goal for the robot is to reach the target without colliding with the obstacle. It should at the same time reach the target as fast as possible. These are conflicting objectives that need to be traded off. In this section, we define the mathematical model of the dynamics and formulate an optimal control problem that will meet the objectives.
II-A System Dynamics
Let $r_{k}=\begin{bmatrix}r_{k}^{x}&r_{k}^{y}\end{bmatrix}^{T}\in\mathbb{R}^{2}$ and $h_{k}=\begin{bmatrix}h_{k}^{x}&h_{k}^{y}\end{bmatrix}^{T}\in\mathbb{R}^{2}$ represent the positions of the robot and the moving obstacle at time $k$, respectively. Moreover, assume that $t\in\mathbb{R}^{2}$ is the target position, where $\mathbb{R}$ is the set of real numbers. The dynamics of the robot and the moving obstacle are described by the following equations:
$\begin{cases}r_{k+1}=r_{k}+u_{k}\triangleq F_{r}\left(r_{k},u_{k}\right)\\ h_{k+1}=h_{k}+w_{k}\triangleq F_{h}\left(h_{k},w_{k}\right)\end{cases}$ (1)
where $F_{r}:\mathbb{R}^{2}\times\mathcal{U}\rightarrow\mathbb{R}^{2}$, $F_{h}:\mathbb{R}^{2}\times\mathcal{W}\rightarrow\mathbb{R}^{2}$. Here $\mathcal{U}=\left\{R_{u}\begin{bmatrix}\cos\left(\alpha)\right)\\ \sin\left(\alpha)\right)\end{bmatrix}\in\mathbb{R}^{2}:R_{u}\in\mathbb{R}_{+},\alpha\in[0,2\pi]\right\}$, and $\mathcal{W}=\left\{R_{w}\begin{bmatrix}\cos\left(\gamma)\right)\\ \sin\left(\gamma)\right)\end{bmatrix}\in\mathbb{R}^{2}:R_{w}\in\mathbb{R}_{+},\gamma\in[0,2\pi]\right\}$. Where $\mathbb{R}_{+}$ is the set of nonnegative real numbers. We let $w$ take different values within the set $\mathcal{W}$ with different probabilities. The models defined in (1) are known as constant position models [14].
Assumption 1
We assume that the robot and the moving obstacle are solid bodies with radius of $\frac{R}{2}$. This assumption simplifies the problem, but we can also handle the general case where the robot and obstacle have different sizes. Moreover, when the distance between the robot and the moving obstacle is less than or equal to $R$, i.e., $d\triangleq||r-h||_{2}\leq R$, the robot has collided with the moving obstacle. Additionally, when the distance between the robot and the target is less than or equal to $R$, i.e. $e\triangleq||r-t||_{2}\leq R$, the robot has reached the target.
II-B Stochastic Infinite Horizon Optimal Control Problem
Assume that $r_{0}$ and $h_{0}$ are the initial positions of the robot and the moving obstacle, respectively. The problem is to find an optimal policy that enables the robot to reach the final state, i.e. the target, while trying to avoid colliding with the moving obstacle. The mathematical description of this problem is as follows:
$\operatorname{minimize}\lim_{N\to\infty}\mathbb{E}\left[\sum_{k=0}^{N-1}f(h_{k},r_{k})\right]$ (2)
subject $\text{subject to}\quad\begin{cases}r_{k+1}=F_{r}\left(r_{k},u_{k}\right)\\ h_{k+1}=F_{h}\left(h_{k},w_{k}\right)\end{cases},\quad k\geq 0$
[…]
where $\mathbb{E}[\cdot]$ denotes mathematical expectation, and where $f:\mathbb{R}^{2}\times\mathbb{R}^{2}\rightarrow\mathbb{R}_{+}$ is the so-called incremental cost. We define this function for a given target position $t$ as
$f(h,r)=\begin{cases}0,\quad||h-t||_{2}\leq R\\ \lambda(||h-t||_{2}-1)^{2}+\frac{1}{||h-r||_{2}+\epsilon},\quad||h-t||_{2}>R\end{cases}$ (3)
=$=\begin{cases}0,\quad e\leq R\\ \lambda(e-1)^{2}+\frac{1}{d+\epsilon},\quad e>R\end{cases}\triangleq\bar{f}(d,e)$
where $\bar{f}:\mathbb{R}_{+}\times\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}$, and where $\lambda\in\mathbb{R}_{++}$ is a tuning parameter, and ϵ++italic-$\epsilon\in\mathbb{R}_{++}$ is a small number. Here, $\mathbb{R}_{++}$ is the set of positive real numbers. The motivation behind selecting this function is that when the robot reaches the target, i.e., $e\leq R$, the incurred cost is zero. However, when the robot gets close to the moving obstacle or is far from the target, the incurred cost is high. Note that when the robot reaches the target, then the incurred cost is zero, making the target position $t$ a so-called cost-free state. In other words, once the robot reaches the target, it will stay there without incurring any cost. This cost free state is well-known in the literature of stochastic shortest path problem [15],[16] and is also referred to as the destination [17].
II-C Solution of Optimal Control
In this section, we will solve the infinite horizon optimal control problem (2) using the Bellman equation:
$V(h,r)=\min_{u\in\mathcal{U}}\mathbb{E}\left[f(h,r)+V\left(F_{h}(h,w),F_{r}(r,u)\right)\right]$ (4)
If we find a solution $V$ to this equation, the optimal policy is the minimizing argument in the right-hand side of the equation. Thus the optimal policy is a function of $(h,r)$. Note that, the Bellman equation in (4) must be solved for all $(h,r)$ and for a given target position $t$. This problem leads to the ’curse of dimensionality’ due to the large state space. We can use value iterations [11] to solve the Bellman equation in (4):
$V_{k+1}(h,r)=\min_{u\in\mathcal{U}}\mathbb{E}\left[f(h,r)+V_{k}\left(F_{h}(h,w),F_{r}(r,u)\right)\right]$ (5)
where $V_{0}(h,r)=0,\forall h,r\in\mathbb{R}^{2}$. The limit of this sequence as $k$ goes to infinity satisfies the Bellman equation.
III Geometric Symmetry
In this section we will show that the value function has a symmetry, and we will use this property to reduce the domain of the value function.
Definition 1
Consider the two pairs, $(h_{1},r_{1})$ and $(h_{2},r_{2})$, in Fig. 1. If $d_{1}=d_{2}$, $e_{1}=e_{2}$, and $\theta_{1}=\pm\theta_{2}$, then these pairs are said to be symmetric around $t$
$x$$y$$\mathbf{t}$$\mathbf{r_{1}}$$\mathbf{r_{2}}$$\mathbf{h_{1}}$$\mathbf{h_{2}}$$\mathbf{e_{1}}$$\phi_{1}$$\mathbf{e_{2}}$$\phi_{2}$$\mathbf{d_{1}}$$\mathbf{d_{2}}$$\theta_{2}$$\theta_{1}$
Figure 1: Geometrical representation of position of robot, and moving obstacle for two symmetric pairs $(h_{1},r_{1})$, and $(h_{2},r_{2})$
Consider two symmetric pairs $(h_{1},r_{1})$ and $(h_{2},r_{2})$. Let $u\in\mathcal{U}$, and $w\in\mathcal{W}$. Furthermore, assume that the probability density function of $w$ only depends on $R_{w}$, i.e. it is constant for each value of $\gamma$. In Fig. 2, two symmetric pairs $(h_{1},r_{1})$ and $(h_{2},r_{2})$, along with their possible future values, are depicted for the case where $R_{u}$ and $R_{w}$ have fixed values. Let $\mu(h,r,R_{u},\alpha)=\mathbb{E}\left[f(h,r)+V_{k}\left(F_{h}(h,w),F_{r}(r,u)\right)\right]$, and $\eta(h,r,R_{u})=\min_{\alpha\in[0,2\pi]}\mu(h,r,R_{u},\alpha)$, where $\mu:\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}_{+}\times[0,2\pi]\rightarrow\mathbb{R}$, and $\eta:\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}_{+}\rightarrow\mathbb{R}$. Then we can rewrite (5) as follows
$V_{k+1}(h,r)=\min_{R_{u}\in\mathbb{R}_{+}}\min_{\alpha\in[0,2\pi]}\mathbb{E}\left[f(h,r)+V_{k}\left(F_{h}(h,w),F_{r}(r,u)\right)\right]$ (6)
$=\min_{R_{u}\in\mathbb{R}_{+}}\min_{\alpha\in[0,2\pi]}\mu(h,r,R_{u},\alpha)=\min_{R_{u}\in\mathbb{R}_{+}}\eta(h,r,R_{u})$
For every fixed value of $R_{u}$, and a minimizing $\alpha_{1}$ corresponding to $r_{1}$, there is a minimizing $\alpha_{2}$ corresponding to $r_{2}$ such that $\mu(h_{1},r_{1},R_{u},\alpha_{1})=\mu(h_{2},r_{2},R_{u},\alpha_{2})$. In Fig. 2, the future values of these two points for a fixed value of $R_{u}$ are depicted by $r_{1}^{+}$ and $r_{2}^{+}$, respectively. Moreover, it holds that $\eta(h_{1},r_{1},R_{u})=\eta(h_{2},r_{2},R_{u})$. Then we optimize over $R_{u}$. In this way we have optimality over all $u\in\mathbb{R}^{2}$. Therefore, the minimum value of the right-hand side of (5) is the same for these two pairs, regardless of the values of $\phi_{1}$ and $\phi_{2}$. Hence for (5), we have
$V_{k+1}\left(h_{1},r_{1}\right)=$ (7)
$V_{k+1}\left(t+e_{1}\begin{bmatrix}\cos(\phi_{1})\\ \sin(\phi_{1})\end{bmatrix}+d_{1}\begin{bmatrix}\cos(\phi_{1}+\theta_{1})\\ \sin(\phi_{1}+\theta_{1})\end{bmatrix},t+e_{1}\begin{bmatrix}\cos(\phi_{1})\\ \sin(\phi_{1})\end{bmatrix}\right)=$
$V_{k+1}\left(t+e_{2}\begin{bmatrix}\cos(\phi_{2})\\ \sin(\phi_{2})\end{bmatrix}+d_{2}\begin{bmatrix}\cos(\phi_{2}+\theta_{2})\\ \sin(\phi_{2}+\theta_{2})\end{bmatrix},t+e_{2}\begin{bmatrix}\cos(\phi_{2})\\ \sin(\phi_{2})\end{bmatrix}\right)$
$=V_{k+1}\left(h_{2},r_{2}\right)$
where $e_{1}=e_{2}$, $d_{1}=d_{2}$, $\theta_{1}=\pm\theta_{2}$ and $\phi_{1},\phi_{2}$ are arbitrary. Hence $V_{k+1}$ can be considered to be a function of only $d$, $e$, and $\theta$. Now, we want to show that also the right-hand side of (5) can be expressed in terms of only $d$, $e$, and $\theta$.
$x$$y$$\mathbf{r}_{1}^{+}$$\mathbf{r}_{2}^{+}$$\mathbf{t}$$\mathbf{r_{1}}$$\mathbf{r_{2}}$$\mathbf{h_{1}}$$\mathbf{h_{2}}$$\mathbf{e}$$\phi_{1}$$\mathbf{e}$$\phi_{2}$$\mathbf{d}$$\mathbf{d}$$\theta_{2}$$\theta_{1}$
Figure 2: Geometrical Representation of Position of Robot, Obstacle, and Target for two symmetric pairs with all possible Future Values
Theorem 1
Let $(h,r)$ be a specific pair defined by $(d,e,\theta)$ as in Fig. 3, i.e.,
$\begin{cases}h=r+d\begin{bmatrix}\cos(\theta)&\sin(\theta)\end{bmatrix}^{T}\\ r=t+e\begin{bmatrix}1&0\end{bmatrix}^{T}\end{cases}$ (8)
Then we can express the next values of $\left(d,e,\theta\right)$ as functions of the current values of $\left(d,e,\theta\right)$ and $u$, and $w$. More precisely, we have
$\begin{cases}e_{+}=F_{e}(e,u)\\ \theta_{+}=F_{\theta}(d,e,\theta,u,w)\\ d_{+}=F_{d}(d,\theta,u,w)\end{cases}$ (9)
where $F_{e}:\mathbb{R}_{+}\times\mathcal{U}\rightarrow\mathbb{R}_{+}$, $F_{\theta}:\mathbb{R}_{+}\times\mathbb{R}_{+}\times[0,\pi]\times\mathcal{U}\times\mathcal{W}\rightarrow[0,\pi]$, and $F_{d}:\mathbb{R}_{+}\times[0,\pi]\times\mathcal{U}\times\mathcal{W}\rightarrow\mathbb{R}_{+}$ are functions defined in the appendix.
Proof:
The proof is given in the appendix. ∎
$x$$y$$\mathbf{t}$$\mathbf{r}$$\mathbf{h}$$\mathbf{r}_{+}$$\mathbf{h}_{+}$$\mathbf{e}_{+}$$\mathbf{e}$$\mathbf{d}_{+}$$\mathbf{d}$$\theta$$\theta_{+}$
Figure 3: Geometrical Representation of Position of Robot, Obstacle, and Target with Future Values
We now define
$W_{k}(d,e,\theta)=V_{k}\left(t+e\begin{bmatrix}1\\ 0\end{bmatrix}+d\begin{bmatrix}\cos(\theta)\\ \sin(\theta)\end{bmatrix},t+e\begin{bmatrix}1\\ 0\end{bmatrix}\right)$ (10)
where $W_{k}:\mathbb{R}_{+}\times\mathbb{R}_{+}\times[0,2\pi]\rightarrow\mathbb{R}$. Then we can equivalently write (5) as
$W_{k+1}(d,e,\theta)$ (11)
$=\min_{u\in\mathcal{U}}\mathbb{E}\left[\bar{f}(d,e)+W_{k}\left(F_{d}(d,\theta,u,w),F_{e}(e,u),F_{\theta}(d,e,\theta,u,w)\right)\right]$
The limit of this recursion will define a solution to (4). This limit is denoted by $W(d,e,\theta)$, where $W:\mathbb{R}_{+}\times\mathbb{R}_{+}\times[0,2\pi]\rightarrow\mathbb{R}$.
IV fitted Value Iteration
In this section, we propose to solve the recursion for $W_{k}$ in (11) using fitted value iterations. To use fitted value iterations, we introduce vector-valued function $\varphi:\mathbb{R}_{+}\times\mathbb{R}_{+}\times[0,\pi]\rightarrow\mathbb{R}^{p}$, known as feature vector, and parameters $a_{k}\in\mathbb{R}^{p}$. Our goal is to approximate the function $W_{k}(d,e,\theta)$, with $\tilde{W}:\mathbb{R}_{+}\times\mathbb{R}_{+}\times[0,\pi]\times\mathbb{R}^{p}\rightarrow\mathbb{R}$. One possibility is to use a linear regression model as follows:
$\tilde{W}(d,e,\theta,a_{k})=a_{k}^{T}\varphi(d,e,\theta)$ (12)
We consider a piecewise constant approximation of $W_{k}$ by assuming that the $i$-th component of $\varphi$, $i\in\{1,2,\cdots,p\}$ is defined as
$\varphi_{i}(d,e,\theta)=\begin{cases}1,\quad(d,e,\theta)\in\mathcal{S}_{i}\\ 0,\quad(d,e,\theta)\notin\mathcal{S}_{i}\end{cases}$ (13)
where $\mathcal{S}_{i}$ is a partition of +×+×$\mathbb{R}_{+}\times\mathbb{R}_{+}\times\left[0,\pi\right]$. More precisely, write $\mathcal{D}=\cup_{j=1}^{N_{d}-1}\mathcal{D}_{j}=[d_{1},d_{N_{d}}]$, where $\mathcal{D}_{j}=[d_{j},d_{j+1}]$, such that $\mathcal{D}_{j}\cap\mathcal{D}_{\bar{j}}=\emptyset$, for $j\neq\bar{j}$. Similarly, we define $\mathcal{E}=\cup_{l=1}^{N_{e}-1}\mathcal{E}_{l}$, and $\Theta=\cup_{m=1}^{N_{\theta}-1}\Theta_{m}$. Then we let $\mathcal{S}_{i}=\mathcal{D}_{j}\times\mathcal{E}_{l}\times\Theta_{m}$, where $j\in\left\{1,2,\cdots,N_{d}-1\right\}$, $l\in\left\{1,2,\cdots,N_{e}-1\right\}$, $m\in\left\{1,2,\cdots,N_{\theta}-1\right\}$, and $i=(j-1)(N_{e}-1)(N_{\theta}-1)+(l-1)(N_{\theta}-1)+m$. This results in $p=(N_{d}-1)(N_{e}-1)(N_{\theta}-1)$ partitions.
We also restrict ourselves to the case that $R_{u}\in\{0,\bar{R}_{u}\}$, and $R_{w}\in\{0,\bar{R}_{w}\}$ to have a tractable problem. Our previous results apply also to this case. We also discretize $\alpha$ and $\gamma$, and define $\bar{\mathcal{U}}=\left\{\bar{R}_{u}\begin{bmatrix}\cos\left(\alpha)\right)\\ \sin\left(\alpha)\right)\end{bmatrix}:\alpha\in\left\{0,\frac{\pi}{n_{1}},\frac{2\pi}{n_{1}},\cdots,\frac{(2n_{1}-1)\pi}{n_{1}}\right\}\right\}\cup\left\{\begin{bmatrix}0\\ 0\end{bmatrix}\right\}$, $\bar{\mathcal{W}}=\left\{\bar{R}_{w}\begin{bmatrix}\cos\left(\gamma)\right)\\ \sin\left(\gamma)\right)\end{bmatrix}:\gamma\in\left\{0,\frac{\pi}{n_{2}},\frac{2\pi}{n_{2}},\cdots,\frac{(2n_{2}-1)\pi}{n_{2}}\right\}\right\}\cup\left\{\begin{bmatrix}0\\ 0\end{bmatrix}\right\}$, where $n_{1}$ and $n_{2}$ are two natural numbers. We can rewrite the set $\bar{\mathcal{W}}$ as $\{w^{0},w^{1},\cdots,w^{2n_{2}}\}$, in which $w^{i}$ is $\begin{bmatrix}\cos\left(\frac{i\pi}{n_{2}}\right)\\ \sin\left(\frac{i\pi}{n_{2}}\right)\end{bmatrix}$, for $10\leq i\leq 2n_{2}-$ 1, and $w^{i}$ is $\begin{bmatrix}0\\ 0\end{bmatrix}$ for $i=2n_{2}$. This discretization results in that our previous results only apply approximately. We refer the reader to [11] for more details about fitted value iterations.
We consider $N_{r}$ different samples $(d^{s},e^{s},\theta^{s})\in\mathbb{R}_{+}\times\mathbb{R}_{+}\times[0,\pi]$ with at least one sample in each set of the partition. Then, at each iteration, we perform two updates:
  1. 1.
    Value Update:
    $\beta_{k}^{s}=\min_{u}\mathbb{E}\left[\bar{f}(d^{s},e^{s})+\tilde{W}\left(d_{+}^{s},e_{+}^{s},\theta_{+}^{s},a_{k}\right)\right]$ (14)
  2. 2.
    Parameter Update:
    $a_{k+1}=\arg\min_{a}\sum_{s=1}^{N_{r}}\left(a^{T}\varphi(d^{s},e^{s},\theta^{s})-\beta_{k}^{s}\right)^{2}$ (15)
where $d_{+}^{s}$, $e_{+}^{s}$, and $\theta_{+}^{s}$ are defined as in (9). We iterate until $\delta_{k}\triangleq||a_{k+1}-a_{k}||_{\infty}$ becomes sufficiently small.
Algorithm 1 summarizes the fitted value iterations.
Algorithm 1 Fitted Value Iteration
1:Input: Incremental Cost $\bar{f}$, Functions $F_{d}$, $F_{e}$ and $F_{\theta}$, Integer $n_{2}$, Control Set $\bar{\mathcal{U}}$, Set $\bar{\mathcal{W}}$, Tolerance $\varepsilon>0$, $N_{r}$ Different Sample Points $(d^{s},e^{s},\theta^{s})$
2:Output: $a$
3:Initialize: $k=0$ ,$a_{0}=0$, Generate $N_{r}$ Sample Points
4:while $||a_{k+1}-a_{k}||_{\infty}>\varepsilon$ do
5:     Value Update:
6:     for $s\leftarrow 0$ to $N_{r}$ do
7:         $Q(u)\leftarrow 0$
8:         for $u\in\bar{\mathcal{U}}$ do
9:              for $i\leftarrow 0$ to $2n_{2}$ do
10:                  Set: $w=w^{i}$
11:                  Compute: $d_{+},e_{+},\theta_{+}$
12:                  Compute:
13:                  $\tilde{W}(d_{+}^{s},e_{+}^{s},\theta_{+}^{s},a_{k})=a_{k}^{T}\varphi(d_{+}^{s},e_{+}^{s},\theta_{+}^{s})$
14:                  Update:
15:                  $Q(u)\leftarrow Q(u)+\frac{1}{2n_{2}+1}\tilde{W}(d_{+}^{s},e_{+}^{s},\theta_{+}^{s},a_{k})$
16:              end for
17:              $Q(u)\leftarrow Q(u)+\bar{f}(d^{s},e^{s})$
18:         end for
19:         $\beta_{k}^{s}=\min_{u\in\bar{\mathcal{U}}}Q(u)$
20:     end for
21:     Parameter Update:
22:     $a_{k+1}=\arg\min_{a}\sum_{s=1}^{N_{r}}\left(a^{T}\varphi(d^{s},e^{s},\theta^{s})-\beta_{k}^{s}\right)^{2}$
23:     $k\leftarrow k+1$
24:end while
V Effect of Probability Distribution and Constraints
In the proposed method, a crucial assumption that needs to be satisfied is the conditions regarding the probability distribution of $w$. In this section, we relax this condition and also consider constraints on the states. As discussed in Section II, if $V$ satisfies the Ballman equation in (4), then minimizing the right-hand side of (4) gives the optimal policy for the specific probability distribution of $w$. We use the optimal value function $V$ that satisfies the Bellman equation in (4) and formulate the following problem:
$u_{k}^{\star}=\operatorname{argmin}_{u_{k}}\mathbb{E}\left[f(h_{k},r_{k})+V(h_{k+1},r_{k+1})\right]$ (16)
subject $\text{subject to}\quad\begin{cases}r_{k+1}=F_{r}\left(r_{k},u_{k}\right)\\ h_{k+1}=F_{h}\left(h_{k},w_{k}\right)\end{cases},\quad k\geq 0$
[…]
where $\mathcal{X}$ is the set of all constraints on $x_{k}$. The expectation is with respect to $w$, and the probability distribution of $w$ is arbitrary. At each time step, we solve the optimization problem (16) in real-time to find the control signal $u_{k}^{\star}$, and then we apply this control signal. Had there been no constraints on $x_{k}$ and the probability distribution had satisfied our previous assumptions, then this online procedure would provide the optimal feedback. However, when we have constraints on the states and the probability distribution does not satisfy our previous assumptions, we only obtain a sub-optimal solution. This approach reduces the impact of the curse of dimensionality at the price of a suboptimal solution.
VI Numerical Results
In this section, we consider a numerical example and we use the proposed method described in Algorithm 1 to find the value function. We let $N_{d}=95$, $N_{e}=65$, $N_{\theta}=26$ , and $\{d_{1},d_{2},\cdots,d_{61}\}=\{0,0.05,0.1,\cdots,3\}$, $\{d_{62},d_{63},\cdots,d_{95}\}=\{3.6,4.4,5.2,\cdots,30\}$. Also we consider $\{e_{1},e_{2},\cdots,e_{31}\}=\{0,0.1,0.2,\cdots,3\}$, $\{e_{32},e_{33},\cdots,e_{65}\}=\{3.6,4.4,5.2,\cdots,30\}$. Moreover we let $\{\theta_{1},\theta_{2},\cdots,\theta_{26}\}=\{0,\frac{\pi}{25},\frac{2\pi}{25},\cdots,\pi\}$. This results in $p=150,400$ partitions. We consider $N_{r}=300,800$ sample points, i.e., two sample points per partition. Additionally, we let $\lambda=0.001$, and $R=1$. We let $h_{0}=\begin{bmatrix}2\\ 6\end{bmatrix}$, $r_{0}=\begin{bmatrix}4\\ 12\end{bmatrix}$, and $t=\begin{bmatrix}4\\ 3\end{bmatrix}$. We consider a probability distribution for $w$ such that its expected value drifts upward and to the right. The sets $\bar{\mathcal{U}}$ and $\bar{\mathcal{W}}$ are defined in Section IV, where $\bar{R}_{u}=\bar{R}_{w}=1$, and we let $n_{1}=n_{2}=16$. Moreover, we define the set $\mathcal{X}$ as follows
$\mathcal{X}=\left\{\begin{bmatrix}h^{x}\\ h^{y}\\ r^{x}\\ r^{y}\end{bmatrix}\in\mathbb{R}^{4}:0\leq r^{x},r^{y},h^{x},h^{y}\leq 20\right\}$
We consider $10000$ different realizations of $w$. We compare the results of the proposed method with receding horizon $A^{\star}$. Table I shows the cost, percentage of collisions, and execution time per stage for these methods, averaged over all different realizations of $w$. The proposed method has a lower cost than receding horizon $A^{\star}$. Additionally, the average computation time per stage and percentage of collisions for the proposed method are less than those for receding horizon $A^{\star}$. We plot one of the samples, and Fig. 4 shows the trajectory of the robot and the moving obstacle for the different methods. From this figure, the trajectory of the robot when using the proposed approach makes a detour to avoid the moving obstacle. Fig. 5 depicts the control signals for the different methods. In this figure, the first component of $u$, denoted by $u^{x}$ is shown in the top plot, and the second component of $u$, denoted by $u^{y}$ is shown in the bottom plot for these two cases. From this figure, the optimal control signal takes different angles, to avoid colliding with the moving obstacle. However, for almost all time instances, the control signal of the $A^{\star}$ approach remains constant, resulting in an almost straight line. Fig. 6 shows the distances between the robot and the target, and between the robot and the moving obstacle for the different methods at each time instance. From this figure, the robot reaches the target faster in the $A^{\star}$ approach than in our proposed method. However, the distance between the robot and the moving obstacle is larger in our proposed method compared to the $A^{\star}$ approach for most of the time instances.
TABLE I: Comparison of proposed method with receding horizon $A^{\star}$
Method Cost Computation Time Per Stage [sec] Percentage of Collision
Proposed Method $3.1168$ $0.0107$ $0.01$
Receding Horizon $A^{\star}$ $4.1683$ $0.2183$ $46.07$
Figure 4: Trajectory of Robot, and Moving Obstacle using Different Methods
Figure 5: Control Signal for different methods
Figure 6: Distance between Robot and Target, and Distance between Robot and Moving obstacle
VII Conclusions
In this paper, we introduced a method to compute an optimal path for a robot in an environment with a stochastically moving obstacle. We reduced the dimension of the domain of the value function. In this way, we have less problem with the curse of dimensionality. When we compare our method with $A^{\star}$, our proposed method outperforms $A^{\star}$. More precisely, our method has lower cost, less computational time per stage, and fewer collisions.
APPENDIX
In this appendix we prove Theorem 1. Let $(h,r)$ be a specific point defined in (8). From Fig. 3, and since from (1) we know that $r_{+}=F_{r}(r,u)$ and $h_{+}=F_{h}(h,w)$, we have
$e_{+}=||r_{+}-t||_{2}=||r+u-t||_{2}$ (17)
$=||t+e\begin{bmatrix}1&0\end{bmatrix}^{T}+u-t||_{2}=||e\begin{bmatrix}1&0\end{bmatrix}^{T}+u||_{2}$
$\triangleq F_{e}(e,u)$
where $F_{e}:\mathbb{R}_{+}\times\mathcal{U}\rightarrow\mathbb{R}_{+}$. To define $\theta_{+}$, let $\nu=r_{+}-t$ and $\xi=h_{+}-r_{+}$, then
$\theta_{+}=\arccos\left(\frac{\nu^{T}\xi}{||\nu||_{2}||\xi||_{2}}\right)\triangleq F_{\theta}(d,e,\theta,u,w)$ (18)
where $F_{\theta}:\mathbb{R}_{+}\times\mathbb{R}_{+}\times[0,\pi]\times\mathcal{U}\times\mathcal{W}\rightarrow[0:\pi]$, and where
$\nu=\begin{bmatrix}e+u^{x}\\ u^{y}\end{bmatrix},\quad\xi=\begin{bmatrix}d\cos(\theta)+w^{x}-u^{x}\\ d\sin(\theta)+w^{y}-u^{y}\end{bmatrix}$ (19)
This means that $\theta_{+}$ is a function of $d$, $e$, $\theta$, $u$, and $w$. Note that in the above equations, $u=\begin{bmatrix}u^{x}\\ u^{y}\end{bmatrix}$, and $w=\begin{bmatrix}w^{x}\\ w^{y}\end{bmatrix}$. Moreover, $d_{+}$ is defined as
$d_{+}=||h_{+}-r_{+}||_{2}$ (20)
$=||d\begin{bmatrix}\cos(\theta)&\sin(\theta)\end{bmatrix}^{T}+w-u||_{2}$
$\triangleq F_{d}(d,\theta,u,w)$
where $F_{d}:\mathbb{R}_{+}\times[0,\pi]\times\mathcal{U}\times\mathcal{W}\rightarrow\mathbb{R}_{+}$. Hence, from (17), (18), and (20), for a specific pair $(h,r)$ defined as (8), $d_{+}$, $e_{+}$, and $\theta_{+}$, can be expressed as functions of $d$, $e$, $\theta$, $u$, and $w$.
ACKNOWLEDGMENT
This work was partially supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation.